# MARS: a Multimodal Alignment and Ranking System for Few-Shot Segmentation

### v1

@@ Abstract
Current Few Shot Segmentation literature lacks a mask selection method that goes beyond visual similarity between the query and example images, leading to suboptimal predictions. We present MARS, a plug-and-play ranking system that leverages multimodal cues to filter and merge mask proposals robustly. Starting from a set of mask predictions for a single query image, we score, filter, and merge them to improve results. Proposals are evaluated using multimodal scores computed at local and global levels. Extensive experiments on COCO-20i, Pascal-5i, LVIS-92i, and FSS-1000 demonstrate that integrating all four scoring components is crucial for robust ranking, validating our contribution. As MARS can be effortlessly integrated with various mask proposal systems, we deploy it across a wide range of top-performer methods and achieve new state-of-the-art results on multiple existing benchmarks. Code will be available upon acceptance.
@@ 1 Introduction
Extending the capabilities of traditional segmentation models to new classes often requires retraining with large annotated datasets. This data-collecting process becomes especially prohibitive in domains such as medical imaging, industrial applications, or satellite imagery. Few Shot Segmentation (FSS) [21, 20] offers a promising alternative by reusing knowledge acquired from diverse classes and tasks, enabling models to adapt to new classes with only a handful of annotated examples. Despite the recent advancements in the field, top-performing methods often fail because of visual dissimilarity between the provided example and the query image.
We contribute to this research field by presenting a Multimodal Alignment and Ranking System (MARS). The method is capable of scoring, filtering, and merging mask proposals from any existing FSS method, effectively boosting the performance of several algorithms across different benchmark datasets. A high-level depiction of the model is presented in Figure1. We achieve this goal leveraging four distinct scores: i) Local Conceptual Score ($LC$), ii) Global Conceptual Score ($GC$), iii) Local Visual Score ($LV$), and iv) Global Visual Score ($GV$). Visual scores are derived from direct correspondences between the support and query images, following established approaches [13, 29]. However, we conjecture that visual cues alone may prove insufficient in challenging scenarios characterized by intraclass variation, occlusions, challenging camera angles, and cut-offs. To overcome these limitations, we introduce two conceptual scores based on the inferred class name and definition of the subject of interest, able to establish a higher-order semantic similarity. Summarizing, our work makes the following contributions:
@@
  • Scoring System: We propose a plug-and-play ranking system introducing three novel scores.
@@ 1.1 Related Works
In contrast, recent developments leverage foundational models like the Segment Anything Model (SAM) [9]. SAM is capable of segmenting any object when provided with various types of prompts (e.g.,, points, coarse masks, bounding boxes), yet it does not inherently offer semantic labeling. Consequently, adapting SAM for FSS shifts the emphasis from direct feature matching to an effective prompting strategy. This often requires identifying candidate points by comparing features from the support examples to those in the query images, using these points to generate prompts that guide SAM in producing candidate mask proposals [30, 23, 13, 27, 29].
Several innovative models have emerged along these lines. For instance, PerSAM [30] introduces a training-free personalization approach for SAM. It uses one-shot data—a reference image and a rough mask of the desired concept—to generate a location confidence map that identifies the target object. Based on the confidence scores, two points are selected as positive and negative priors, which are then encoded as prompt tokens for SAM’s decoder to produce the segmentation.
@@
Figure 2: Overview of the proposed MARS model. The diagram illustrates the key components and interactions within the framework. We highlight the Textual Information Retrieval, Visual-Text Alignment, AlphaClip, Visual-Visual Alignment, and Filtering-Merging Modules in different colors.
@@ 2 Method
    Visual Visual Alignment Module Embedding features from the query image $I_{q}$ and support set $S(l)$ are extracted using a ViT Model [18]. These features produce two key outputs: a Refined Visual Alignment and a Cost Matrix, which contribute to computing the LV and GV scores, respectively.
@@
A complete overview of our pipeline is shown in Fig. 2, with complete details provided in the following sections.
@@ 2.1 Textual Information Retrieval Module
We employ a pre-trained ViP-LLaVA to extract the name of the entity of interest from the support image, i.e., the class name. For this purpose, the image is visually prompted with a red mask contour around the target object and is presented at a 50% zoom level, ensuring a balance between emphasizing object details and preserving contextual information. In $N$-shot settings, where the support set consists of $N$ image-mask pairs, ViP-LLaVA is queried for each image individually, and a majority voting scheme is applied to select the most frequent class name for subsequent processing. This solution enables us to quickly extend all the modules based on textual information to the N-shot scenario.
In addition to the class name, we derive a broader textual description of the entity by leveraging WordNet. The extracted class name is used to query WordNet, which returns a set of candidate synsets. If a single synset is returned, its definition is directly adopted as the object’s description. In cases where multiple synsets are retrieved, indicating potential ambiguity, we prompt ViP-LLaVA (using the same configuration as for class name extraction) to generate a description of the entity. The VLM-generated description is then compared with the candidate definitions from WordNet, and the definition with the maximum words overlap is selected as the final class description. If no overlap is found between the VLM-generated description and any of the WordNet definitions, an empty description is used. Note that the description produced by ViP-LLaVA is used solely for matching purposes, as it may incorporate visual context specific to the support image.
The extracted textual information is used in two ways in the remainder of the pipeline. The Visual-Text Alignment Extraction module employs the class name to determine local conceptual information that enables the computation of the Local Conceptual Score (LC). The AlphaClip Module uses the derived description together with the class name to produce the Global Conceptual Score (GC).
@@ 2.2 Visual-Text Alignment Module
This module generates a Refined Text Alignment, a saliency map that highlights regions in the query image where the subject described by the class name is likely located. Leveraging the visual-text alignment capabilities of a pre-trained CLIP model [19], the module first produces a coarse Text Alignment. Then it refines it through the PI-CLIP [26] Prior Information Refinement (PIR) process followed by min-max normalization. The final Refined Text Alignment is later used to compute the Local Conceptual Score by being multiplied element-wise with each binary mask proposal, with the resulting non-zero values averaged to yield a single score for ranking.
@@
To enhance spatial coherence and semantic precision of the text prior, we employ the Prior Information Refinement (PIR) module [26], which aggregates and normalizes the self-attention map $A$ from the CLIP vision encoder. A binary box mask $B$ is then derived from $T_{A}$ via thresholding (with a threshold of 0.4, following [26, 12]). The Refined Text Alignment ($RTA$) is computed as:
@@ 2.4 Visual-Visual Alignment Module
The design of the two scores ensures that both global alignment and localized feature correspondences are effectively captured, thereby complementing the textual cues provided by the other modules in our pipeline.
@@ 3.3 A Simple MARS Pipeline
In this experiment, we assess the autonomous potential of MARS in a single/few-shot segmentation pipeline that relies on minimal prior information. To isolate the effectiveness of our ranking system, we developed a simplified pipeline based on the SAM Automatic Mask Generator (AMG). In this setup, SAM generates mask proposals from a uniform random sampling grid of points—without any guidance regarding the target object. The main motivation of this experiment is to evaluate how well MARS can perform when the quality of mask proposals is deliberately suboptimal, thus removing any bias introduced by advanced FSS-specific proposal methods. This allows us to assess the intrinsic strength of our ranking and merging strategy in handling noisy or imprecise proposals.
Table 4 provides a comparative analysis between the proposed AMG + MARS pipeline and several published methods on the one-shot segmentation task. Despite the simplicity of the mask proposal stage, our approach surpasses specialist models (e.g., HSNet [16], VAT [8]) and remains competitive with top-performing methods (e.g., SegGPT, Matcher, GF-SAM). Although the overall performance is lower than that achieved by methods with more sophisticated proposal mechanisms, these results underscore the strenght of our ranking system and highlight the importance of an efficient mask proposal stage in FSS.
Table 4: Performance of well-known methods on one-shot and five-shot segmentation tasks. The table reports mIoU scores.

### v2

@@ Abstract
Few Shot Segmentation aims to segment novel object classes given only a handful of labeled examples, enabling rapid adaptation with minimal supervision. Current literature crucially lacks a selection method that goes beyond visual similarity between the query and example images, leading to suboptimal predictions. We present MARS, a plug-and-play ranking system that leverages multimodal cues to filter and merge mask proposals robustly. Starting from a set of mask predictions for a single query image, we score, filter, and merge them to improve results. Proposals are evaluated using multimodal scores computed at local and global levels. Extensive experiments on COCO-20i, Pascal-5i, LVIS-92i, and FSS-1000 demonstrate that integrating all four scoring components is crucial for robust ranking, validating our contribution. As MARS can be effortlessly integrated with various mask proposal systems, we deploy it across a wide range of top-performer methods and achieve new state-of-the-art results on multiple existing benchmarks. Code will be available upon acceptance.
@@ 1 Introduction
Extending the capabilities of traditional segmentation models to new classes often requires retraining with large annotated datasets. This data-collecting process becomes especially prohibitive in domains such as medical imaging, industrial applications, or satellite imagery. Few Shot Segmentation (FSS) [21, 20] offers a promising alternative by reusing knowledge acquired from diverse classes and tasks, enabling models to adapt to new classes with only a handful of annotated examples. Despite the recent advancements in the field, top-performing methods often fail because of visual dissimilarity between the provided example and the query image. We contribute to this research field by presenting a Multimodal Alignment and Ranking System (MARS). The method is capable of scoring, filtering, and merging mask proposals from any existing FSS method, effectively boosting the performance of several algorithms across different benchmark datasets. A high-level depiction of the model is presented in Figure 1. We achieve this goal leveraging four distinct scores: i) Local Conceptual Score ($LC$), ii) Global Conceptual Score ($GC$), iii) Local Visual Score ($LV$), and iv) Global Visual Score ($GV$). Visual scores are derived from direct correspondences between the support and query images, following established approaches [13, 29]. However, we conjecture that visual cues alone may prove insufficient in challenging scenarios characterized by intraclass variation, occlusions, challenging camera angles, and cut-offs. To overcome these limitations, we introduce two conceptual scores based on the inferred class name and definition of the subject of interest, able to establish a higher-order semantic similarity. Summarizing, our work makes the following contributions:
@@
  • Scoring System: We propose a plug-and-play ranking with four scores: three novel and one EMD based.
@@ 2 Related Works
In contrast, recent developments leverage foundational models like the Segment Anything Model (SAM) [9]. SAM is capable of segmenting any object when provided with various types of prompts (e.g. , points, coarse masks, bounding boxes), yet it does not inherently offer semantic labeling. Consequently, adapting SAM for FSS shifts the emphasis from direct feature matching to an effective prompting strategy. This often requires identifying candidate points by comparing features from the support examples to those in the query images, using these points to generate prompts that guide SAM in producing candidate mask proposals [30, 23, 13, 27, 29].
Several innovative models have emerged along these lines. For instance, PerSAM [30] introduces a training-free personalization approach for SAM. It uses one-shot data, a reference image and a rough mask of the desired concept, to generate a location confidence map that identifies the target object. Based on the confidence scores, two points are selected as positive and negative priors, which are then encoded as prompt tokens for SAM’s decoder to produce the segmentation.
@@ 3 Method
    Visual-Visual Alignment Module Embedding features from the query image $I_{q}$ and support set $S(l)$ are extracted using a ViT Model [18]. These features produce two key outputs: a Refined Visual Alignment and a Cost Matrix, which contribute to computing the LV and GV scores, respectively.
@@
Figure 2: Overview of the proposed MARS model. The diagram illustrates the key components and interactions within the framework. We highlight the Textual Information Retrieval, Visual-Text Alignment, Visual-Visual Alignment, and the Score-Filter-Merge Modules in different colors.
A complete overview of our pipeline is shown in Fig. 2, with complete details provided in the following sections and Fig. 3.
@@ 3.1 Textual Information Retrieval Module
We employ pretrained ViP-LLaVA to extract the class name from the support image by overlaying a red contour on the target object and applying a 50% zoom (as detailed in Section  6.2 of Supplemental Material), balancing emphasis on object detail with contextual information. In $N$-shot settings, where the support set consists of $N$ image-mask pairs, ViP-LLaVA is queried for each image individually, and a majority voting scheme is applied to select the most frequent class name for subsequent processing. This solution enables us to quickly extend all the modules based on textual information to the N-shot scenario.
In addition to the class name, we derive a broader textual description of the entity by leveraging WordNet. The extracted class name is used to query WordNet, which returns a set of candidate synsets. If a single synset is returned, its definition is directly adopted as the object’s description. In cases where multiple synsets are retrieved, indicating potential ambiguity, we prompt ViP-LLaVA (using the same configuration as for class name extraction) to generate a description of the entity.
Figure 3: Detailed view of the Visual-Text and Visual-Visual Alignment modules. The Visual-Text Alignment module (top) extracts features from the query image using CLIP and aligns them with foreground and background textual features to generate the Softmax-GradCAM map, referred to as the unrefined text prior. This prior is subsequently refined through the PIR module, adapted from [26]. The Visual-Visual Alignment module (bottom) computes similarity between query and support features using cosine similarity. Foreground and background priors are obtained by aggregating the respective similarity matrices, and the resulting visual prior is refined using the PIR module for downstream processing.
The VLM-generated description is then compared with the candidate definitions from WordNet, and the definition with the maximum words overlap is selected as the final class description. If no overlap is found between the VLM-generated description and any of the WordNet definitions, an empty description is used. Note that the description produced by ViP-LLaVA is used solely for matching purposes, as it may incorporate visual context specific to the support image.
The extracted textual information is used in two ways in the remainder of the pipeline. The Visual-Text Alignment module employs the class name to determine local conceptual information that enables the computation of the Local Conceptual Score (LC). The AlphaClip Module uses the derived description together with the class name to produce the Global Conceptual Score (GC).
@@ 3.2 Visual-Text Alignment Module
This module generates a Refined Text Alignment, a saliency map that highlights regions in the query image where the subject described by the class name is likely located. Leveraging the visual-text alignment capabilities of a pre-trained CLIP model [19], the module first produces a coarse Text Alignment. Then it refines it through the PI-CLIP [26] Prior Information Refinement (PIR) process followed by min-max normalization. The final Refined Text Alignment is later used to compute the Local Conceptual Score by being multiplied element-wise with each binary mask proposal, with the resulting non-zero values averaged to yield a single score for ranking. A depiction of the model is provided in the upper part of Fig. 3.
@@
To enhance spatial coherence and semantic precision of the text prior, we employ the Prior Information Refinement (PIR) module [26], which aggregates and normalizes the self-attention map $A$ from the CLIP vision encoder. A binary box mask $B$ is then derived from $TA$ via thresholding (with a threshold of 0.4, following [26, 12]). The Refined Text Alignment ($RTA$) is computed as:
@@ 3.4 Visual-Visual Alignment Module
The design of the two scores ensures that both global alignment and localized feature correspondences are effectively captured, thereby complementing the textual cues provided by the other modules in our pipeline. A depiction of the model is provided in the bottom part of Fig. 3.
@@ 4.3 A Simple MARS Pipeline
In this scenario, we assess the autonomous potential of MARS in a single/few-shot segmentation pipeline that relies on minimal prior information. To isolate the effectiveness of our ranking system, we developed a simplified pipeline based on the SAM Automatic Mask Generator (AMG). In this setup, SAM generates mask proposals from a uniform random sampling grid of points, without any guidance regarding the target object. Further information about the SAM mask generation procedure used can be found in the Supplementary Materials. The main motivation of this experiment is to evaluate how well MARS can perform when the quality of mask proposals is deliberately suboptimal, thus removing any bias introduced by advanced FSS-specific proposal methods. This allows us to assess the intrinsic strength of our ranking and merging strategy in handling noisy or imprecise proposals.
Table 4 provides a comparative analysis between the proposed AMG + MARS pipeline and most relevant published methods. Despite the simplicity of the mask proposal stage, our approach has the highest score on COCO-20i (1-shot) and remains competitive with top-performing methods in other settings. These results underscore the strength of our ranking system, asserting the dominance of the mask selection procedure over the mask proposal stage in FSS.
Table 4: Performance of well-known methods on one-shot and five-shot segmentation tasks. The table reports mIoU scores. We have highlighted in bold the top performer method and underlined the second-best.